# Key-Point-Driven Mathematical Reasoning Distillation of Large Language Model

### v3

@@ Abstract
Large Language Models (LLMs) have demonstrated exceptional proficiency in mathematical reasoning tasks due to their extensive parameter counts and training on vast datasets. Despite these capabilities, deploying LLMs is hindered by their computational demands. Distilling LLM mathematical reasoning into Smaller Language Models (SLMs) has emerged as a solution to this challenge, although these smaller models often suffer from errors in calculation and semantic understanding. Prior work has proposed Program-of-Thought Distillation (PoTD) to avoid calculation error. To further address semantic understanding errors, we propose Key-Point-Driven Mathematical Reasoning Distillation (KPDD). KPDD enhances the reasoning performance of SLMs by breaking down the problem-solving process into three stages: Core Question Extraction, Problem-Solving Information Extraction, and Step-by-Step Solution. This method is further divided into KPDD-CoT, which generates Chain-of-Thought rationales, and KPDD-PoT, which creates Program-of-Thought rationales. The experiment results show that KPDD-CoT significantly improves reasoning abilities, while KPDD-PoT achieves state-of-the-art performance in mathematical reasoning tasks. Our approach effectively mitigates misunderstanding errors, advancing the deployment of efficient and capable SLMs.
@@
Xunyu Zhu1,2, Jian Li1,2thanks: Corresponding author, Can Ma1,2, Weiping Wang1,2 1Institute of Information Engineering, Chinese Academy of Sciences 2School of Cyber Security, University of Chinese Academy of Sciences {zhuxunyu, lijian9026, macan, wangweiping}@iie.ac.cn
@@ 1 Introduction
Large language models (LLMs) have achieved impressive performance in mathematical reasoning tasks. Recent work (Wang et al., 2023) further proposes Chain-of-Thought (CoT) to enhance the mathematical reasoning abilities of LLMs. However, the massive scale of LLMs presents significant challenges for deployment.
@@
Figure 1: Error analysis of 50 GSM8K problems with incorrect answers returned by CoTD using FlanT5-Base. The experimental results indicate that multiple errors may exist in the reasoning process of CoTD, with understanding errors and calculation errors being the major factors affecting CoTD’s reasoning performance.
@@
A feasible solution to address this problem is to use black-box distillation to transfer mathematical reasoning abilities from LLMs to SLMs. Chain-of-Thought Distillation (CoTD) Ho et al. (2023) is a representative mathematical reasoning distillation method. CoTD prompts LLMs to generate reasoning processes for each question, constructing a mathematical reasoning dataset. This dataset is then used to fine-tune SLMs, enhancing their mathematical reasoning abilities. However, there remains a significant performance gap between SLMs and LLMs. Prior work (Wei et al., 2022) identifies three main error types in CoT reasoning: Calculation errors, Missing Step errors, and Semantic misunderstanding errors. To explore the reasons for the performance gap between SLMs and LLMs, we conducted the same error analysis on CoTD. Our preliminary experiments (shown in Figure 1) reveal numerous error combinations in CoTD, with calculation and semantic misunderstanding errors being the most prevalent. Prior work Zhu et al. (2023) proposed Program-of-Thought Distillation (PoTD) to avoid calculation errors by formulating the reasoning process as a Python program executed by an external interpreter. This approach allows the SLM to focus on generating the program, avoiding calculation errors and improving reasoning performance. Given these circumstances, our paper focuses on addressing semantic misunderstanding errors in CoTD to further enhance the reasoning performance of SLMs.
@@
In our paper, we propose a novel mathematical reasoning distillation method called Key-Point-Driven Mathematical Reasoning Distillation (KPDD) to enhance the mathematical reasoning performance of SLMs. KPDD breaks the reasoning process into three parts: (1) Core Question Extraction: Identifies the core question from the original problem. (2) Problem-Solving Information Extraction: Extracts relevant data and information needed to solve the problem. (3) Step-by-Step Solution: Uses the extracted key points to solve the problem in a step-by-step manner. The third part is further divided into two formats, KPDD-CoT and KPDD-PoT: (1) KPDD-CoT: Generates rationales in the form of Chain-of-Thought (CoT). This method focuses on reducing misunderstanding errors and explicitly illustrates the reasoning process, aiding in error analysis. (2) KPDD-PoT: Generates rationales in the form of Program-of-Thought (PoT). This approach not only reduces misunderstanding errors but also avoids calculation errors, further enhancing the SLM’s mathematical reasoning performance.
@@ 2.1 Chain-of-Thought Reasoning
Chain-of-Thought refers to prompt LLMs to solve the question step by step. Prior work Wei et al. (2022) find that Chain-of-Thought can effectively improve the reasoning performance of LLMs. Based on the findings, Kojima et al. Kojima et al. (2022) further introduce zero-shot CoT, which significantly improves the model’s reasoning performance by simply adding the prompt "Let’s think step by step" before answering. To avoid calculation error in CoT, Chen et al. Chen et al. (2023) formulate the reasoning process into program. Furthermore, Wang et al. Wang et al. (2023) introduce a self-consistency decoding strategy, which generates diverse reasoning paths and then selects the most consistent answer by considering these paths comprehensively. Least-to-most prompting Zhou et al. (2023) breaks down complex problems into a series of simpler subproblems and solves these subproblems sequentially. Zhong et al. Zhong et al. (2024) encourage LLMs to deeply understand problems and leverage key information for better reasoning. Inspired by these methods, our work introduces Key-Point-Driven Mathematical Reasoning Distillation (KPDD) to further enhance SLMs’ mathematical reasoning.
@@ 2.2 Black-box Distillation
Currently, close-source LLMs usually have stronger reasoning performance than open-source LLMs. However, in general, we can only obtain the output of close-source LLMs. Based on this situation, black-box distillation is proposed to distill abilities from LLMs to SLMs. Specifically, black-box distillation first prompts LLMs to generate a distillation dataset, and then this dataset is used to fine-tune SLMs to improve their reasoning performance. For example, Ho et al. Ho et al. (2023) prompt LLMs to generate a CoT reasoning dataset, which was then used to fine-tune an SLM, thereby enhancing its reasoning ability. Shridhar et al. Shridhar et al. (2023) utilize a LLM to train a problem decomposer and a subproblem solver separately. They used the decomposer to break down the original problem into multiple subproblems, and then employed the subproblem solver to solve each subproblem individually. By integrating the solutions to these subproblems, they derived the final answer. Fu et al. Fu et al. (2023) analyze CoT distillation and find that there is a trade-off between the specific abilities and the general abilities of SLMs. When CoT distillation is used to enhance the specific abilities of SLMs, their general abilities decrease correspondingly. Zhu et al. Zhu et al. (2023) construct a distillation dataset, where the reasoning format changed from CoT reasoning to PoT reasoning. Through this method, the SLM can delegate the calculation process to an additional Python interpreter to avoid calculation errors, thereby enhancing the SLM’s mathematical reasoning performance. Zhu et al. Zhu et al. (2024) further formalize the reasoning process into equations and find that a diverse range of reasoning formats can effectively enhance the mathematical reasoning performance of SLMs. Our work introduces a novel distillation approach where two SLMs independently extract the core question and key problem-solving information from an original question. These key points are then utilized to guide another SLM in solving the original question effectively.
@@ 3 Method
Figure 2: Detailed data generation for KPDD-CoT. Initially, we use few-shot prompting to guide LLMs in producing reasoning processes. We then extract the correct answers from these processes, discarding any reasoning process that does not align with the correct answers. In this way, we have constructed a high-quality KPDD-CoT reasoning distillation dataset.
In this work, we introduce a novel distillation method for mathematical reasoning tasks called Key-Point-Driven Distillation (KPDD), structured into three stages: (1) Stage 1: KPDD distills the first SLM to extract the core question from the original question. (2) Stage 2: KPDD distills the second SLM to extract problem-solving information from the original question. (3) Stage 3: KPDD distills the third SLM to solve the original problem using the core question and problem-solving information. In Stage 3, we prompt the LLM to construct two types of reasoning datasets: (1) CoT Rationales: These are more comprehensible to both humans and LLMs, showcasing a detailed reasoning process. (2) PoT Rationales: These rationales delegate computational tasks to an external Python interpreter, thereby avoiding calculation errors.
@@ 3.1 Data Generation from LLMs
When given a mathematical dataset, our primary task is to utilize an LLM to generate a reasoning process for each mathematical problem in the dataset. By doing so, we augment the mathematical dataset to construct the distillation dataset. Furthermore, in stage 3, our KPDD method employs two distillation approaches: one distills the SLM to generate CoT rationales for problem-solving, and the other distills the SLM to generate PoT rationales for problem-solving. In other words, our KPDD method can be divided into two approaches: KPDD-CoT and KPDD-PoT.
3.1.1 Data Generation for KPDD-CoT
Our method uses few-shot prompting to prompt the LLM to synthesize the reasoning dataset. Figure 2 reveals the distillation dataset generation process. Specifically, we first randomly sample $k$ data pairs $(x,y)$ from a mathematical dataset $\mathcal{D}$, where $x$ represents the question and $y$ represents the answer. Then, for these sampled data, we manually construct the reasoning process $c$. Each reasoning process $c$ includes a core question, problem-solving information, and rationales in CoT format. These elements are separated by HTML tags: "<core>{core question}</core><info>{problem-solving information}</info><cot>{rationales in CoT format}</cot>". Finally, we obtain a demonstration dataset $\mathcal{D}_{c}$. At the same time, we further introduce an instruction "Firstly, let’s extract the most comprehensive and detailed key question. Then, let’s identify and list the most useful information related to the question. Finally, let’s understand the key question and the problem-solving information, solve the question step by step, and show the answer.". By leveraging the demonstration set and the instruction, we prompt the LLM to generate the reasoning process for the mathematical questions. The KPDD-CoT dataset generation is formalized as:
@@ 3.1.1 Data Generation for KPDD-CoT
Figure 3: Detailed data generation for KPDD-PoT. Similar to KPDD-CoT, we prompt LLMs to generate the reasoning process in the KPDD-PoT format. We extract the rationale in the PoT format from this process and run it through a Python interpreter. If there are errors or incorrect answers, we discard the reasoning process. Finally, we constructed a high-quality KPDD-PoT reasoning dataset.
3.1.2 Data Generation for KPDD-PoT
Similar to KPDD-CoT, we first sample $k$ examples from a mathematical dataset $\mathcal{D}$, then use these samples to manually create a demonstration set $\mathcal{D}_{p}$. Each reasoning process in the demonstration set $\mathcal{D}_{p}$ includes a core question, problem-solving information, and rationales in PoT format. These elements are also separated by HTML tags: "<core>{core question}</core><info>{problem-solving information}</info><pot>{rationales in PoT format}</pot>". Then, we further introduce the instruction a instruction "Firstly, let’s extract the most comprehensive and detailed key question. Then, let’s identify and list the most useful information related to the question. Finally, let’s understand the key question and the problem-solving information, and generate the python code (return ans) to solve the question.". By utilizing the demonstration set $\mathcal{D}_{p}$ and the instruction, we prompt LLMs to generate reasoning processes in for questions. Figure 3 shows the data synthesis process for KPDD-PoT, and the KPDD-PoT dataset generation process can be formalized as:
$p_{i}=f_{\mathcal{M}}(x_{i},\mathcal{D}_{p}).$ (2)
@@ 3.2 Fine-tuning SLMs
After constructing these reasoning datasets, we use them to fine-tune the SLMs. In the KPDD, we fine-tune three SLMs: the first SLM, called KPDD-CoT/PoT-core, is used to extract the core question from the original problem, the second SLM, called KPDD-CoT/PoT-info, extracts the problem-solving information, and the third SLM, called KPDD-CoT/PoT-solve, uses both the core question and problem-solving information to solve the original question.
3.2.1 Fine-tuning SLMs for KPDD-CoT
Firstly, we construct a core question subset from the KPDD-CoT dataset, denoted as $\mathcal{D}_{CC}$. Each sample in this subset can be represented as $(x,cc)$, where $x$ represents the original question and $cc$ represents the core question. For each training instance $(x,cc)$ from $\mathcal{D}_{CC}$, we prepend the prompt $p_{cc}$Let’s extract the most comprehensive and detailed core question." to the question $x$. This guides the KPDD-CoT-core in fine-tuning to accurately extract the corresponding core question $cc$. The fine-tuning loss function can be represented as follows:
$\mathcal{L}=-\sum_{i=1}^{N}\sum_{t=1}^{T}\log P({cc}^{i}_{t}\mid{cc}^{i}_{<t},x^{i},p_{cc}),$ (3)
where $N$ is the number of examples in $\mathcal{D}_{CC}$, $p_{cc}$ is the prompt, and ${cc}_{:T}$ is the sequence of the core question.
Then, we construct a problem-solving subset from the KPDD-CoT dataset, denoted as $\mathcal{D}_{CI}$. Each sample in this subset can be represented as $(x,ci)$, where $x$ represents the original question and $ci$ represents the problem-solving information. For each data point $(x,ci)$ from $\mathcal{D}_{CI}$, we add the prompt $p_{ci}$Let’s identify and list the most useful information related to the question." to the question $x$. This guides the KPDD-CoT-info in fine-tuning to accurately extract the corresponding problem-solving information $ci$. The fine-tuning loss function can be represented as follows:
$\mathcal{L}=-\sum_{i=1}^{N}\sum_{t=1}^{T}\log P({ci}^{i}_{t}\mid{ci}^{i}_{<t},x^{i},p_{ci}),$ (4)
where $N$ is the number of examples in $\mathcal{D}_{CI}$, $p_{ci}$ is the prompt, and ${ci}_{:T}$ is the sequence of the problem-solving information.
Figure 4: The inference process of KPDD. When given an original question, KPDD-CoT/PoT-core and KPDD-CoT/PoT-info first extract the core question and the problem-solving information. Then, KPDD-CoT/PoT-solve uses these key points to generate rationales to solve the original question.
Finally, we construct a problem-solving subset from the KPDD-CoT dataset, denoted as $\mathcal{D}_{CS}$. Each sample in this subset can be represented as $(x,cc,ci,cs)$, where $x$ represents the original question, $cc$ represents the core question, $ci$ represents the problem-solving information, and $cs$ represents the rationales in CoT format. For each data $(x,cc,ci,cs)$ from $\mathcal{D}_{CS}$, we integrate the original question $x$, the core question $cc$, the problem-solving information $ci$ and the prompt $p_{cs}$Let’s understand the core question and the problem-solving information, solve the question step by step, and show the answer." to construct a new input. This guides the KPDD-CoT-solve in fine-tuning to generate rationales $cs$ for solving the origin question in CoT format. The fine-tuning loss function can be represented as follows:
$\mathcal{L}=-\sum_{i=1}^{N}\sum_{t=1}^{T}\log P({cs}^{i}_{t}\mid{cs}^{i}_{<t},x^{i},cc^{i},ci^{i},p_{cs}),$ (5)
@@ 3.2.1 Fine-tuning SLMs for KPDD-CoT
3.2.2 Fine-tuning SLMs for KPDD-PoT
In KPDD-PoT, aside from replacing the KPDD-CoT dataset with the KPDD-PoT dataset, the fine-tuning method for KPDD-PoT-core remains consistent with that of KPDD-CoT-core, and the fine-tuning method for KPDD-PoT-info remains consistent with that of KPDD-CoT-info. However, the fine-tuning method of KPDD-PoT-solve is different with KPDD-CoT-solve. The main difference between them is the input instruction. Specifically, when fine-tuning KPDD-PoT-solve, the input instruction is: "Let’s understand the core question and the problem-solving information, and generate the python code (return ans) to solve the question." This instruction guides the model to not only understand the core question and the problem-solving information but also to generate Python code that can compute the answer. This approach leverages the model’s ability to perform code generation, which can be particularly effective for solving mathematical problems programmatically.

### v4

@@ Abstract
Large Language Models (LLMs) excel in mathematical reasoning due to their extensive parameters and training data, but their high computational demands hinder deployment. Distilling LLM reasoning into Smaller Language Models (SLMs, $\leq$ 1B parameters) is a potential solution, yet these models often struggle with calculation and semantic errors. Previous work introduced Program-of-Thought Distillation (PoTD) to reduce calculation mistakes. To tackle semantic errors, we propose Key-Point-Driven Mathematical Reasoning Distillation (KPDD), which improves SLM reasoning by splitting the problem-solving process into Key Points Extraction and Step-by-Step Solution. KPDD includes KPDD-CoT, generating Chain-of-Thought rationales, and KPDD-PoT, producing Program-of-Thought rationales. Experiments show KPDD-CoT enhances reasoning capabilities, while KPDD-PoT achieves state-of-the-art performance in mathematical tasks, effectively reducing misunderstanding errors and promoting the deployment of efficient, capable SLMs.
@@
Xunyu Zhu1,2, Jian Li3thanks: Corresponding author, Can Ma1,2, Weiping Wang1,2 1Institute of Information Engineering, Chinese Academy of Sciences 2School of Cyber Security, University of Chinese Academy of Sciences 3 School of Artificial Intelligence, Beijing Normal University {zhuxunyu, macan, wangweiping}@iie.ac.cn, jli@bnu.edu.cn
@@ 1 Introduction
Large language models (LLMs) OpenAI (2023); Yang et al. (2024); Bi et al. (2024); Jiang et al. (2024); Zeng et al. (2024); Cai et al. (2024) have achieved impressive performance in mathematical reasoning tasks. Recent work Wang et al. (2023) further proposes Chain-of-Thought (CoT) to enhance the mathematical reasoning abilities of LLMs. However, the massive scale of LLMs presents significant challenges for deployment.
@@
Figure 1: Error analysis of 50 GSM8K problems with incorrect answers returned by CoTD using FlanT5-Base. The experimental results indicate that understanding errors and calculation errors are the major factors affecting CoTD’s reasoning performance.
@@
A feasible solution to address this problem is to use black-box distillation to transfer mathematical reasoning abilities from LLMs to SLMs. Chain-of-Thought Distillation (CoTD) Ho et al. (2023) is a representative mathematical reasoning distillation method. CoTD prompts LLMs to generate reasoning processes for each question, constructing a mathematical reasoning dataset. This dataset is then used to fine-tune SLMs, enhancing their mathematical reasoning abilities. However, there remains a significant performance gap between SLMs and LLMs. Prior work Wei et al. (2022) identifies three main error types in CoT reasoning: Calculation errors, Missing Step errors, and Semantic misunderstanding errors. To explore the reasons for the performance gap between SLMs and LLMs, we conducted the same error analysis on CoTD. Our preliminary experiments (shown in Figure 1) indicate that the most common issues affecting CoTD’s mathematical reasoning performance are calculation errors and semantic misunderstandings. Prior work Zhu et al. (2023) proposed Program-of-Thought Distillation (PoTD) to avoid calculation errors by formulating the reasoning process as a Python program executed by an external interpreter. This approach allows the SLM to focus on generating the program, avoiding calculation errors and improving reasoning performance. Given these circumstances, our paper focuses on addressing semantic misunderstanding errors in CoTD to further enhance the reasoning performance of SLMs.
@@
In our paper, we propose a novel mathematical reasoning distillation method called Key-Point-Driven Mathematical Reasoning Distillation (KPDD) to enhance the mathematical reasoning performance of SLMs. KPDD breaks the reasoning process into two parts: (1) Key points Extraction: identifies and extracts the core question and the relevant information from the original problem. (2) Step-by-Step Solution: uses the extracted key points to solve the problem in a step-by-step manner. The second part is further divided into two formats, KPDD-CoT and KPDD-PoT: (1) KPDD-CoT: Generates rationales in the form of Chain-of-Thought (CoT). This method focuses on reducing misunderstanding errors and explicitly illustrates the reasoning process, aiding in error analysis. (2) KPDD-PoT: Generates rationales in the form of Program-of-Thought (PoT). This approach not only reduces misunderstanding errors but also avoids calculation errors, further enhancing the SLM’s mathematical reasoning performance.
@@ 2.1 Chain-of-Thought Reasoning
Chain-of-Thought refers to prompt LLMs to solve the question step by step. Prior work Wei et al. (2022) finds that Chain-of-Thought can effectively improve the reasoning performance of LLMs. Based on the findings,  Kojima et al. (2022) further introduce zero-shot CoT, which significantly improves the model’s reasoning performance by simply adding the prompt "Let’s think step by step" before answering. To avoid calculation error in CoT,  Chen et al. (2023) formulate the reasoning process into program.  Wang et al. (2023) introduce a self-consistency decoding strategy, which generates diverse reasoning paths and then selects the most consistent answer by considering these paths comprehensively. Least-to-most prompting Zhou et al. (2023) breaks down complex problems into a series of simpler subproblems and solves these subproblems sequentially.  Zhong et al. (2024) encourage LLMs to deeply understand problems and leverage useful information for better reasoning. Inspired by these methods, our work introduces Key-Point-Driven Mathematical Reasoning Distillation (KPDD), which generate distillation dataset with abundant mathematical reasoning knowledge to enhance SLMs’ mathematical reasoning.
@@ 2.2 Black-box Distillation
Currently, LLMs usually have stronger reasoning performance than SLMs ($\leq$ 1B). However, in general, we can only obtain the output of close-source LLMs. Based on this situation, black-box distillation is proposed to distill abilities from LLMs to SLMs. Specifically, black-box distillation first prompts LLMs to generate a distillation dataset, and then this dataset is used to fine-tune SLMs to improve their reasoning performance. For example,  Ho et al. (2023) prompt LLMs to generate a CoT reasoning dataset, which was then used to fine-tune an SLM, thereby enhancing its reasoning ability.  Hsieh et al. (2023) introduce a multi-task distillation framework that enables the SLM to focus not only on predicting the final output labels but also on predicting intermediate reasoning steps.  Shridhar et al. (2023) utilize a LLM to train a problem decomposer and subproblem solver, breaking problems into subproblems and solving them individually to get the final answer.  Fu et al. (2023) analyze CoT distillation, finding a trade-off: enhancing specific SLM abilities decreases their general abilities.  Zhu et al. (2023) create a PoT reasoning dataset, allowing SLMs to use Python interpreters for better mathematical reasoning and error avoidance.  Zhu et al. (2024) further formalize the reasoning process into equations and find that a diverse range of reasoning formats can effectively enhance the mathematical reasoning performance of SLMs. In contrast to the methods mentioned above, our work presents a novel distillation approach in which one SLM is employed to extract the core question and problem-solving information from the original question. These key points are then used to guide a second SLM in effectively solving the original question.
@@ 3 Method
Figure 2: Detailed data generation for KPDD-CoT. Initially, we use few-shot prompting to guide LLMs in producing reasoning processes. We then discard any reasoning process that does not align with the correct answers. In this way, we have constructed a high-quality KPDD-CoT reasoning distillation dataset.
In this work, we introduce a novel distillation method for mathematical reasoning tasks called Key-Point-Driven Distillation (KPDD), structured into two stages: (1) Stage 1: KPDD distills the first SLM to extract the core question and the problem-solving information from the original question. (2) Stage 2: KPDD distills the second SLM to solve the original problem using the core question and problem-solving information.
@@ 3.1 Data Generation from LLMs
When given a mathematical dataset, our primary task is to utilize an LLM to generate a reasoning process for each mathematical problem in the dataset. By doing so, we augment the mathematical dataset to construct the distillation dataset. Furthermore, in stage 2, our KPDD method employs two distillation approaches: one distills the SLM to generate CoT rationales for problem-solving, and the other distills the SLM to generate PoT rationales for problem-solving. In other words, our KPDD method can be divided into two approaches: KPDD-CoT and KPDD-PoT. Here is a data generation process for KPDD-CoT. Furthermore, the data generation process for KPDD-PoT is similar to that of KPDD-CoT, with the detailed procedure for KPDD-PoT provided in Appendix A.
Our method uses few-shot prompting to prompt the LLM to synthesize the reasoning dataset. Figure 2 reveals the distillation dataset generation process. Specifically, we first randomly sample $k$ data pairs $(x,y)$ from a mathematical dataset $\mathcal{D}$, where $x$ represents the question and $y$ represents the answer. Then, for these sampled data, we manually construct the reasoning process $c$. Each reasoning process $c$ includes a core question, problem-solving information, and rationales in CoT format. Finally, we obtain a demonstration dataset $\mathcal{D}_{c}$. At the same time, we further introduce an instruction "Firstly, let’s extract the most comprehensive and detailed core question. Then, let’s identify and list the most useful information related to the question. Finally, let’s understand the core question and the problem-solving information, solve the question step by step, and show the answer.". By leveraging the demonstration set and the instruction, we prompt the LLM to generate the reasoning process for the mathematical questions. The KPDD-CoT dataset generation is formalized as:
@@ 3.2 Fine-tuning SLMs
After constructing these reasoning datasets, we use them to fine-tune the SLMs. In the KPDD, we fine-tune two SLMs: the first SLM, called KPDD-CoT/PoT-Key, is used to extract the key points from the original problem, the second SLM, called KPDD-CoT/PoT-Solve, solve the original question based on these key points. Here is a fine-tuning process for KPDD-CoT. The fine-tuning process for KPDD-PoT is similar to that of KPDD-CoT, with the detailed procedure for KPDD-PoT provided in Appendix B.
Firstly, we construct a key points subset from the KPDD-CoT dataset, denoted as $\mathcal{D}_{KP}$. Each sample in this subset can be represented as $(x,kp)$, where $x$ represents the original question and $kp$ represents key points, which consist of the core question and problem-solving information. For each training instance $(x,kp)$ from $\mathcal{D}_{KP}$, we prepend the prompt $p_{kp}$Let’s extract the core question and the most useful information from the question." to the question $x$. This guides the KPDD-CoT-Key in fine-tuning to effectively extract the corresponding key points $kp$. The fine-tuning loss function can be represented as follows:
$\mathcal{L}=-\sum_{i=1}^{N}\sum_{t=1}^{T}\log P({kp}^{i}_{t}\mid{kp}^{i}_{<t},x^{i},p_{kp}),$ (2)
where $N$ is the number of examples in $\mathcal{D}_{KP}$, $p_{kp}$ is the prompt, and ${kp}_{:T}$ is the sequence of the key points.
Then, we construct a problem-solving subset from the KPDD-CoT dataset, denoted as $\mathcal{D}_{CS}$. Each sample in this subset can be represented as a tuple $(x,kp,cs)$, where $x$ represents the original question, $kp$ represents the key points, encompassing the core question and problem-solving information, and $cs$ represents the rationales in CoT format. For each data from $\mathcal{D}_{CS}$, we integrate the original question, the key points and the prompt $p_{cs}$Let’s understand the core question and the problem-solving information, solve the question step by step, and show the answer." to construct a new input. This guides the KPDD-CoT-Solve in fine-tuning to generate rationales $cs$ for solving the origin question in CoT format. The fine-tuning loss function can be represented as follows:
$\mathcal{L}=-\sum_{i=1}^{N}\sum_{t=1}^{T}\log P({cs}^{i}_{t}\mid{cs}^{i}_{<t},x^{i},kp^{i},p_{cs}),$ (3)